OpenStudy (anonymous):

A doctor prescribes an ointment that is 4% hydrocortisone. A pharmacist has 1% and 5% concentrations in stock. How many ounces of each should the pharmacist use to make a 6-ounce tube?

OpenStudy (anonymous):

\[.05x+.01(6-x)=.04\times 6\] solve for \(x\)

OpenStudy (anonymous):

probably easier to multiply both sides by 100 to get rid of annoying decimal and solve \[5x+(6-x)=24\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw btw i hope it is clear that \(x\) is the number of ounces of 5% cream

OpenStudy (anonymous):

yes I understand thank you